OpenStudy (anonymous):

If y varies directly as x and y = 6 when x = 18, then find x when y = 10.

OpenStudy (campbell_st):

ok... so direct variation means the equation is in the form y = kx you need to find k, knowing when x = 18, y = 6 so whats the value of k...?

OpenStudy (campbell_st):

and can you please check the question says y = 6 when x = 18... and not the other way round.

OpenStudy (anonymous):

1/3? @campbell_st

OpenStudy (campbell_st):

thats correct so the equation is y = 1/3 x so now subtitute y = 10 to find x

OpenStudy (anonymous):

So the answer is 10/3?

OpenStudy (campbell_st):

no its 10 = 1/3 x so try again

OpenStudy (campbell_st):

just multiply both sides of the equation to find x

OpenStudy (anonymous):

30!

OpenStudy (campbell_st):

yes... well done

OpenStudy (anonymous):

@campbell_st Thank you so much!

OpenStudy (anonymous):

@campbell_st I have one more question. Question: Table represents direct variation x: 1,2,3,4 y: 12,6,4,3 This is false right?

OpenStudy (campbell_st):

ok... so again y = kx use the 1st result to find k 12 = k* 1 so k = 12 then the equation is y = 12x is that true for every pair try x = 2 what should you get and compare it to the y value in the table then x = 3 what should you get

OpenStudy (anonymous):

It is false ?

OpenStudy (campbell_st):

opps... yes you are correct

OpenStudy (anonymous):

Okay! Thank you very much, again!